# Quantum enhanced real-time sensing of protein-gold adsorption kinetics

### v2

@@ Abstract
Analyzing the kinetics of biological processes plays a significant role in understanding fundamental cellular functions. Many physics-based technologies used to study such processes are limited by the shot noise inherent to the coherent states of light. These technologies can greatly benefit by leveraging quantum probes to improve the sensitivity of measurements in cellular biology. Surface Plasmon Resonance (SPR) technique has been used effectively to achieve label-free, real-time measurements of protein binding kinetics, which constitutes an important biological phenomenon occurring near the cell membrane. Here, we demonstrate the integration of this technique with the two-mode bright squeezed state having fewer fluctuations as compared to the coherent state to improve the sensitivity of measurement in studying a protein-gold adsorption process. We show 4dB of squeezing as we record the signal-to-noise ratio as the function of time and it is maintained throughout the kinetic process. The quantum advantage as shown in terms of squeezing is achieved despite the total absorption of from the source until the final detection after the sensor. Overall, we provide the most practical setup for improving the sensitivity of the time-dependent measurements involved in various biological processes at the molecular level.
@@ I Introduction
However, the Surface Plasmon Resonance (SPR) has proved to be a powerful technique in medical research mainly used to analyze events at or near a metal-dielectric interface, where the metal is typically gold film or nanoparticlesand the dielectric constitutes the analyte under study [11, 12, 13]. It has been used effectively to achieve label-free, real-time measurements of protein binding kinetics which constitutes an important dynamic process occurring near the cell membrane [14, 15, 16]. The SPR technique is based on sensing the minute changes in refractive index of the sample which in turn gives information about concentration, affinity and the kinetic parameters of the protein binding event. Understanding the dynamic processes such as protein binding is essential in learning how diseases develop at the molecular level and finding the potential treatment strategies [17]. Protein binding plays an important role in the area of drug design and delivery and studying this process requires more sophisticated tools to advance the field of medical research [18, 19, 20, 21].
@@
In our work we consider the Two Mode Bright Squeezed State (TMBSS) because of its continuous wave nature allows for a higher number of photons as compared to the discrete quantum sources while still maintaining the quantum correlations between the two modes. It can enhance the sensing capabilities specifically by reducing the fluctuations associated with the intensity difference signal. Pooser et al. has demonstrated a quantum plasmonic sensor using the Krescthcmann configuration and the TMBSS source to measure the shift in SPR dip for samples with different refractive index [31]. However, no previous study has experimentally investigated the quantum advantage of TMBSS in the context of time dependent SPR sensing to measure the protein-gold adsorption kinetics. Protein adsorption on gold induces a response similar to the protein binding process as it changes the refractive index of the sample when the protein associates with or dissociates from the gold film. Any further references to protein binding to the gold film are written in the context of the adsorption process.
In this paper, we present the most practical setup to integrate the TMBSS along with the SPR sensor and give proof of principle experiment showing the quantum advantage which stays almost constant throughout the protein adsorption process.Here, we mainly focus on the association of the protein with the gold surface. This setup can be used to measure the kinetic parameters with higher sensitivity than its classical counterpart when working in the shot noise limited region. We first explain in detail the workings of the SPR sensor that detects minute changes in the refractive index of the sample in contact with the gold film. The sensitivity of measurement of refractive index is directly related to the sensitivity of measuring the kinetic parameters. We then explain the noise reduction achieved using the TMBSS in the form of squeezing. We describe the experimental setup along with the data processing performed to measure the SNR. We present our findings from plotting the SPR dip useful in deciding the incident angle on sensor to quantifying the SNR as a function of time as the BSA protein binds to the gold film on the sensor. We show improvement in the sensitivity of estimating the observable rate constant of the binding process. Finally we discuss the challenges, the significance as well as the future implications of the setup described in this study.
@@ II Schematics of the quantum enhanced surface plasmon sensor
Surface Plasmon Resonance is a well-known technique in the field of bio sensing. It is based on the phenomenon where the surface plasmon (electron oscillations) at the metal-dielectric interface gets coupled with the evanescent wave generated due to an incident electromagnetic field [32, 33, 34]. This results in the formation of Transverse Magnetic (TM) wave confined to the metal-dielectric interface called as the Surface Plasmon Polariton (SPP) wave. Although the phenomenon could be observed using different coupling arrangements, the most common setup used is the Krescthcmann geometry as shown in Fig. 1(a). Here the prism, the gold film, and the dielectric medium form the apparatus useful in sensing the refractive index of the dielectric. The SPP wave propagates along the x-axis and decays exponentially on both sides of the interface along the z-axis as shown in Fig. 1(a).
@@
Thus, at a particular angle of incidence after the critical angle the SPP are formed leading to the absorption of energy from the incident field. This is called the Surface Plasmon Resonance angle. As we plot the reflectivity of this sensor against the angle of incidence, we can see an absorption dip at the angle of resonance. Thus excitation of the SPP wave is highly sensitive to the refractive index of the dielectric medium at the interface. If we keep the thickness of the metal and the frequency of the laser constant, even the slightest change in the refractive index of the sample can be detected using the shift in the resonance dip. This is shown by the plot in Fig. 1(b) which is based on the theoretical framework called Transfer-matrix method [35] which describes propagation of an electromagnetic wave through a homogeneous film situated between two semi-infinite media as explained in detail in the appendix section. Note that the reflectance curve is more linear on the left side of the resonance angle as compared to the right side. Operating in this region produces higher shift in intensity measurement for a small shift in the resonance angle and avoids saturation of the measurements. As shown by the marker in Fig. 1(b) we operate on the left side of the resonance angle as it allows for a higher dynamic range in measurements accommodating various analyte concentrations without requiring frequent recalibration. This region offers more stability considering the mechanical vibrations or thermal drift allowing for more reproducible and reliable measurements.
@@
where r is the squeezing parameter and the being the Heisenberg operators representing the electromagnetic field propagation as described in Fig. 2(a). Notice that the input modes, the coherent state and the vacuum state are the single modes. Using the balanced detection technique as described in [37], the signal is given by and the associated fluctuations are given by .
@@ III.1 Generating Two-Mode Bright Squeezed State
To generate the Two Mode Bright Squeezed State we use a process called degenerate Four Wave Mixing (FWM) in which annihilation of two photons of same frequency leads to the creation of two new photons of slightly different frequencies. This non-linear process occurs inside the Rb85 atomic vapor contained in a 2.5 cm long cylindrical vapor cell. The atomic ensemble serves as a non-linear medium where the line corresponding to the transition including the hyperfine structure of the Rb85 atom constitutes a double- energy level configuration as shown in Fig. 3(b).
Thus, all the four control parameters namely the single photon detuning, two photon detuning, the pump, probe alignment, and the temperature of the cell are optimized to achieve the squeezing of 7.8 dB as measured at the source just after the cell. Please refer[38, 39, 36, 40] for more details of generating the TMBSS source using FWM in Rb85 atomic vapor. The flip mirrors in Fig. 3(a), create a path for the classical beam which technically is the seed probe diverted just before entering the Rb cell. The further arrangements constitute the application of the twin beams in studying the protein-gold binding process.
@@ III.2 Applying TMBSS to the SPR sensor
Next, the probe beam reaches the sensor used to study the BSA protein adsorption on gold film. The sensor is assembled using commercially available SPR chip along with the on-campus facilities as explained in appendix B. The flow cell cavity as shown in Fig. 3(c) is filled up with the required sample using the tubes connected via a peristaltic pump. In this paper, we use BSA solution which was prepared by dissolving 1gm Bovine Serum Albumin (BSA) Lyophilized powder into 10 ml of deionized water. We wait for approximately one hour to ensure complete dissolution and uniformity of the solution before injecting the fluid into the flow cell. The sensor assembly is then installed on a rotation stage such that the reflectivity can be measured at different incident angles. All the angles are measured using the incident angle of as the reference angle which is experimentally fixed using two irises.
@@ IV Results
Classical results: Firstly, we found the Surface Plasmon Resonance (SPR) angle for deionized water (n=1.33) by plotting the reflectivity of the sensor against the scanning angle (angle made with the normal of the prism-air interface) using a coherent source. The sensor was aligned such that the classical beam in the probe path is incident at zero degrees to the normal of prism-air interface (incident angle ) and reflects at a right angle to the incident beam. We fix this position using two irises on the reflected path as a reference point. We recorded the reflectivity of the sensor which is the ratio of reflected intensity with the incident intensity against the scanning angle at the prism-gold interface as shown in Fig. 6(a). We collected the data in the scanning range of to the normal of the prism-air interface with increments of . This gives the incident angle at the gold-prism interface in the range . We find that the intensity reaches its minimum at an angle of () at which point around of the intensity is absorbed by the excited SPPs. This aligns closely with the theoretical expectations as shown in Fig. 1(b). The deviation from the theoretical simulation can be attributed to experimental errors corresponding to the initial angle. For better comparability the angle on the x-axis was converted into the normalized scale such that the angle of resonance is considered to be at zero. Fig. 6(a) helps decide the position at which we should fix the angle to realize the maximum change in reflected intensity with a minimum shift in the refractive index. The reason to choose the left side of the curve is discussed earlier in the schematics section. We also provide the experimental data for a shift in angle of resonance from BSA (deionized water) to BSA solution along with the theoretical expectation for the resonance angles corresponding to different BSA concentrations in Appendix D.
@@
Squeezed light characterization: Squeezing in dB as defined by Eq. (5) depends on the absorption that changes at different incident angles on the SPR sensor as shown in Fig.7(a). We experimentally collected the noise power at different angles around the absorption dip and compared it with the SNL at each point. The SNL was recorded for one angle and the rest of the points were calculated using the Poisson characteristic of the coherent source . The coherent source is verified to be shot-noise limited around the working frequency of 2MHz as shown in Appendix C. As shown in Fig. 7(a), we get squeezing of around 4 dB at an angle of 0.8 degrees left of as decided using the classical SPR dip data. The squeezing as a function of reflectivity is reported in Fig.7(b).
@@
where is the initial reflected intensity at time , is the maximum change in reflected intensity as it reaches equilibrium and is the observable rate constant in units of . The approximate mean value of the rate constant can be found out using the non-liner fit as shown in Fig. 9. For our setup, we find the values as follows , and . Here we focus mainly on showing the improvement in the sensitivity of estimating the parameter when the signal is shot-noise limited. Eq. (8) can be written as,
@@ V Conclusion
In summary, we have demonstrated the quantum advantage of using Two Mode Bright Squeezed State (TMBSS) in improving the sensitivity of a Surface Plasmon Resonance (SPR) based sensor to study biokinetic processes. As an example of one such process, we used the protein – Bovine Serum Albumin (BSA) binding to the gold film on the surface of the sensor. As the protein binds, the reflectivity of the sensor goes up due to the shift in surface plasmon resonance angle. We showed 4dB of squeezing (difference between the coherent source noise and the twin beam noise) as we recorded the signal-to-noise ratio as the function of time and is maintained throughout the binding process. The quantum advantage as shown in terms of squeezing is achieved despite the total absorption of from the source until the final detection after the sensor.
We successfully demonstrated the reduction in the state noise of the source which contributes to a improvement in the sensitivity of estimation of the binding parameter .Thus, integration of a well stabilized source of TMBSS with SPR can achieve more precise measurements of the rate of protein binding to gold surpassing the limitations of the current shot-noise limited techniques. It can also be used to study protein-protein, protein-ligand biding processes that are very significant in the field of biochemistry and biophysics.
Our work reinforces the utility of TMBSS source highlighting its primary advantage: the continuous variable nature as opposed to the single photon sources that allows for a greater number of photons while still employing the quantum mechanical benefits that arise from the intensity difference squeezing. TMBSS constitutes a readily available quantum probe which is better than the coherent source because of its sub-shot noise characteristic thus allowing for lower intensities with comparable results, making it more gentle to the biological samples. The combination of TMBSS along with the sub-diffraction-limited sensing property of the SPR presents a highly promising approach in the realm of cellular biology.
Overall, this project brings together two state-of-the-art techniques and provides a most practical, quantum-enhanced SPR setup aimed at improving the sensitivity of time-dependent measurements in the study of biological processes.
@@ Acknowledgements.
The authors are grateful for the support of the Air Force Office of Scientific Research (Award No. FA-9550-20-1-0366) and the Robert A. Welch Foundation (A-1943-20210327). The final measurements for the SNR were done in the quantum sensing lab at the University of Tennessee at Chattanooga (UTC). We wish to acknowledge the technical help provided by Charles Wallace at Texas AM University.Appendix A: Transfer-Matrix method
@@
Appendix B: Sensor assembly
@@
Appendix C: Shot Noise Limitation
We verified the coherent source for its shot-noise limit. The two coherent beams were directed in the two detectors A and B replicating the arrangement for TMBSS source. The verification was done using the oscilloscope for voltage readings and the spectrum analyzer was used to record the noise power around the working frequency of 2MHz. As shown in Fig. 11 the noise power plotted on y axis is taken relative to the noise power corresponding to the total intensity of 1.5 V. As we can see the SNL increases by 3dB when the intensity is doubled demonstrating the linear dependence of the variance with respect to the mean of the signal.
Appendix D: SPR characteristics for different concentrations
We recorded the reflectivity of the sensor against the scanning angle for BSA (deionized water) and BSA solution as shown in Fig. 12 using the coherent source. Next, we provide the theoretical expectation for the behavior of the SPR resonance angle for various concentrations of BSA solution as explained below.
We use the information provided in [42] to find the refractive index of various concentrations of the BSA solution as shown in Fig. 13(a). We then map these refractive indices to the angle of resonance calculated using the Transfer Matrix Method (TMM) as explained in appendix A and the schematics section in this paper. Thus we plot the resonance angle as a function of the concentration of BSA in water as shown in Fig. 13(b). The plot shows the linearity as expected for a calibration curve.

### v3

@@ Abstract
Analyzing the kinetics of biological processes plays a significant role in understanding fundamental cellular functions. Many physics-based technologies used to study such processes are limited by the shot noise inherent to the coherent states of light. These technologies can greatly benefit by leveraging quantum probes to improve the sensitivity of measurements in cellular biology. Surface Plasmon Resonance (SPR) technique has been used effectively to achieve label-free, real-time measurements of protein binding kinetics, which constitutes an important biological phenomenon occurring near the cell membrane. Here, we demonstrate the integration of this technique with the two-mode bright squeezed state having fewer fluctuations as compared to the coherent state to improve the sensitivity of measurement in studying a protein-gold adsorption process. We show 4dB of squeezing as we record the signal-to-noise ratio as the function of time and it is maintained throughout the kinetic process. The improved signal-to-noise ratio leads to a improvement in the sensitivity of measuring the observable rate constant . The quantum advantage as shown in terms of squeezing is achieved despite the total absorption of from the source until the final detection after the sensor. Overall, we provide the most practical setup for improving the sensitivity of the time-dependent measurements involved in various biological processes at the molecular level.
@@ I Introduction
However, the Surface Plasmon Resonance (SPR) has proved to be a powerful technique in medical research mainly used to analyze events at or near a metal-dielectric interface, where the metal is typically gold film or nanoparticles and the dielectric constitutes the analyte under study [11, 12, 13]. It has been used effectively to achieve label-free, real-time measurements of protein binding kinetics which constitutes an important dynamic process occurring near the cell membrane [14, 15, 16]. The SPR technique is based on sensing the minute changes in refractive index of the sample which in turn gives information about concentration, affinity and the kinetic parameters of the protein binding event. Understanding the dynamic processes such as protein binding is essential in learning how diseases develop at the molecular level and finding the potential treatment strategies [17]. Protein binding plays an important role in the area of drug design and delivery. Enhancing the sensitivity of the SPR instrument for analyzing protein binding kinetics can influence the efficacy and safety of a drug and can significantly advance the field of medical research [18, 19, 20, 21].
@@
In our work we consider the Two Mode Bright Squeezed State (TMBSS) because of its continuous wave nature allows for a higher number of photons as compared to the discrete quantum sources while still maintaining the quantum correlations between the two modes. It can enhance the sensing capabilities specifically by reducing the fluctuations associated with the intensity difference signal. Pooser et al. has demonstrated a quantum plasmonic sensor using the Kretschmann configuration and the TMBSS source to measure the shift in SPR dip for samples with different refractive index [31]. However, no previous study has experimentally investigated the quantum advantage of TMBSS in the context of time dependent SPR sensing to measure the protein-gold adsorption kinetics. Protein adsorption on gold induces a response similar to the protein binding process as it changes the refractive index of the sample when the protein associates with or dissociates from the gold film. Any further references to protein binding to the gold film are written in the context of the adsorption process.
In this paper, we present the most practical setup to integrate the TMBSS along with the SPR sensor and give proof of principle experiment showing the quantum advantage which stays almost constant throughout the protein adsorption process. Here, we mainly focus on the association of the protein with the gold surface. This setup can be used to measure the kinetic parameters with higher sensitivity than its classical counterpart when working in the shot noise limited region. We first explain in detail the workings of the SPR sensor that detects minute changes in the refractive index of the sample in contact with the gold film. The sensitivity of measurement of refractive index is directly related to the sensitivity of measuring the kinetic parameters. We then explain the noise reduction achieved using the TMBSS in the form of squeezing. We describe the experimental setup along with the data processing performed to measure the SNR. We present our findings from plotting the SPR dip useful in deciding the incident angle on sensor to quantifying the SNR as a function of time as the Bovine Serum Albumin (BSA) protein binds to the gold film on the sensor. We show improvement in the sensitivity of estimating the observable rate constant of the binding process. Finally we discuss the challenges, the significance as well as the future implications of the setup described in this study.
@@ II Schematics of the quantum enhanced surface plasmon sensor
Surface Plasmon Resonance is a well-known technique in the field of bio sensing. It is based on the phenomenon where the surface plasmon (electron oscillations) at the metal-dielectric interface gets coupled with the evanescent wave generated due to an incident electromagnetic field [32, 33, 34]. This results in the formation of Transverse Magnetic (TM) wave confined to the metal-dielectric interface called as the Surface Plasmon Polariton (SPP) wave. Although the phenomenon could be observed using different coupling arrangements, the most common setup used is the Kretschmann geometry as shown in Fig. 1(a). Here the prism, the gold film, and the dielectric medium form the apparatus useful in sensing the refractive index of the dielectric. The SPP wave propagates along the x-axis and decays exponentially on both sides of the interface along the z-axis as shown in Fig. 1(a).
@@
Thus, at a particular angle of incidence after the critical angle the SPP are formed leading to the absorption of energy from the incident field. This is called the Surface Plasmon Resonance angle. As we plot the reflectivity of this sensor against the angle of incidence, we can see an absorption dip at the angle of resonance. Thus excitation of the SPP wave is highly sensitive to the refractive index of the dielectric medium at the interface. If we keep the thickness of the metal and the frequency of the laser constant, even the slightest change in the refractive index of the sample can be detected using the shift in the resonance dip. This is shown by the plot in Fig. 1(b) which is based on the theoretical framework called Transfer-matrix method [35] which describes propagation of an electromagnetic wave through a homogeneous film situated between two semi-infinite media as explained in detail in the appendix A. Note that the reflectance curve is more linear on the left side of the resonance angle as compared to the right side. Operating in this region produces higher shift in intensity measurement for a small shift in the resonance angle and avoids saturation of the measurements. As shown by the marker in Fig. 1(b) we operate on the left side of the resonance angle as it allows for a higher dynamic range in measurements accommodating various analyte concentrations without requiring frequent recalibration. This region offers more stability considering the mechanical vibrations or thermal drift allowing for more reproducible and reliable measurements.
@@
where is the squeezing parameter and are the Heisenberg operators representing the electromagnetic field propagation as described in Fig. 2(a). This parameter is proportional to the pump power and the third order nonlinear susceptibility which describes the four wave mixing process. The pump is responsible for the FWM process. Notice that the input modes, the coherent state and the vacuum state are the single modes. Using the balanced detection technique as described in [37], the signal is given by and the associated fluctuations are given by .
@@ III.1 Generating Two-Mode Bright Squeezed State
To generate the Two Mode Bright Squeezed State we use a process called degenerate Four Wave Mixing (FWM) in which annihilation of two photons of same frequency leads to the creation of two new photons of slightly different frequencies. This non-linear process occurs inside the Rb85 atomic vapor contained in a 2.5 cm long cylindrical vapor cell. The FWM process is explained in detail in appendix B.
As shown in the upper half of Fig. 3(a) two beams namely the pump and the seed probe interact with the Rb-85 atoms generating two quantum-correlated beams called the probe and the conjugate. In this process, four control parameters - ,, pump-probe alignment, and temperature of the cell are optimized to achieve the squeezing of 7.8 dB as measured at the source just after the cell. Please refer[38, 39, 36, 40] for more details of generating the TMBSS source using FWM in Rb85 atomic vapor. The flip mirrors in Fig. 3(a), create a path for the classical beam which technically is the seed probe diverted just before entering the Rb cell. The further arrangements constitute the application of the twin beams to study the protein-gold binding process.
@@ III.2 Applying TMBSS to the SPR sensor
Next, the probe beam reaches the sensor used to study the BSA protein adsorption on gold film. The sensor is assembled using commercially available SPR chip along with the on-campus facilities as explained in appendix C. The flow cell cavity as shown in Fig. 3(c) is filled up with the required sample using the tubes connected via a peristaltic pump. In this paper, we use BSA solution which was prepared by dissolving 1gm Bovine Serum Albumin (BSA) Lyophilized powder into 10 ml of deionized water. We wait for approximately one hour to ensure complete dissolution and uniformity of the solution before injecting the fluid into the flow cell. The sensor assembly is then installed on a rotation stage such that the reflectivity can be measured at different incident angles. All the angles are measured using the incident angle of as the reference angle which is experimentally fixed using two irises.
@@ IV Results
Classical results: Firstly, we found the Surface Plasmon Resonance (SPR) angle for deionized water (n=1.33) by plotting the reflectivity of the sensor against the scanning angle (angle made with the normal of the prism-air interface) using a coherent source. The sensor was aligned such that the classical beam in the probe path is incident at zero degrees to the normal of prism-air interface (incident angle ) and reflects at a right angle to the incident beam. We fix this position using two irises on the reflected path as a reference point. We recorded the reflectivity of the sensor which is the ratio of reflected intensity with the incident intensity against the scanning angle at the prism-gold interface as shown in Fig. 6(a). We collected the data in the scanning range of to the normal of the prism-air interface with increments of . This gives the incident angle at the gold-prism interface in the range . We find that the intensity reaches its minimum at an angle of () at which point around of the intensity is absorbed by the excited SPPs. This aligns closely with the theoretical expectations as shown in Fig. 1(b). The deviation from the theoretical simulation can be attributed to experimental errors corresponding to the initial angle. For better comparability the angle on the x-axis was converted into the normalized scale such that the angle of resonance is considered to be at zero. Fig. 6(a) helps decide the position at which we should fix the angle to realize the maximum change in reflected intensity with a minimum shift in the refractive index. The reason to choose the left side of the curve is discussed earlier in the schematics section. We also provide the experimental data for a shift in angle of resonance from BSA (deionized water) to BSA solution along with the theoretical expectation for the resonance angles corresponding to different BSA concentrations in Appendix E.
@@
Squeezed light characterization: Squeezing in dB as defined by Eq. (5) depends on the absorption that changes at different incident angles on the SPR sensor as shown in Fig.7(a). We experimentally collected the noise power at different angles around the absorption dip and compared it with the SNL at each point. The SNL was recorded for one angle and the rest of the points were calculated using the Poisson characteristic of the coherent source . The coherent source is verified to be shot-noise limited around the working frequency of 2MHz as shown in Appendix D. As shown in Fig. 7(a), we get squeezing of around 4 dB at an angle of 0.8 degrees left of as decided using the classical SPR dip data. The squeezing as a function of reflectivity is reported in Fig.7(b).
@@
where is the initial reflected intensity at time , is the maximum change in reflected intensity as it reaches equilibrium and is the observable rate constant in units of . The approximate mean value of the rate constant can be found out using the non-liner fit as shown in Fig. 9. For our setup, we find the values as follows , and . The statistical errors in these parameters are expected to be few percent based on the error bars shown in Fig. 9. These error bars were obtained by taking 10 shots for each time t. Here we focus mainly on showing the improvement in the sensitivity of estimating the parameter when the signal is shot-noise limited. Eq. (7) can be written as,
@@
In protein(Ligand)-protein(Analyte) type of interaction the rate constants are defined as , where is the ligand concentration and , are the association and dissociation rate constants respectively. As clearly explained in Eq.(1.13) in [41] analyzing the in combination with the dissociation rate analysis after the equilibration time can be used to find the binding affinity . This parameter or its reciprocal are direct indicators of how strongly the analyte (for eg. drug protein) binds to the ligand (target protein). This is an invaluable source of information in field of drug design and discovery, and improving the sensitivity of measuring these parameters can greatly enhance drug’s efficacy and safety.
@@ V Conclusion
In summary, we have demonstrated the quantum advantage of using Two Mode Bright Squeezed State (TMBSS) in improving the sensitivity of a Surface Plasmon Resonance (SPR) based sensor to study biokinetic processes. As an example of one such process, we used the protein – Bovine Serum Albumin (BSA) binding to the gold film on the surface of the sensor. We showed 4dB of squeezing (difference between the coherent source noise and the twin beam noise) as we recorded the signal-to-noise ratio as the function of time and is maintained throughout the binding process. The quantum advantage as shown in terms of squeezing is achieved despite the total absorption of from the source until the final detection after the sensor. We successfully demonstrated how reduction in the state noise of the source contributes to a improvement in the sensitivity of estimation of the observable rate constant which is a crucial parameter in finding the affinity of a protein binding to another protein.
Our work also reinforces the utility of TMBSS source highlighting its primary advantage: the continuous variable nature as opposed to the single photon sources that allows for a greater number of photons while still employing the quantum mechanical benefits that arise from the intensity difference squeezing. TMBSS constitutes a readily available quantum probe which is better than the coherent source because of its sub-shot noise characteristic thus allowing for lower intensities with comparable results, making it more gentle to the biological samples. The combination of TMBSS along with the sub-diffraction-limited sensing property of the SPR presents a highly promising approach in the realm of molecular and cellular biology.
Thus, integration of a well stabilized source of TMBSS with SPR can achieve more precise measurements of protein-gold binding rate surpassing the limitations of the current shot-noise limited techniques. This improvement in the sensitivity as presented in our study can be easily extended to the kinetic analysis of the protein-protein (ligand-analyte) binding processes that are very significant in the field of biochemistry especially in drug design and discovery. Overall, this project brings together two state-of-the-art techniques and provides a most practical, quantum-enhanced SPR setup aimed at improving the sensitivity of time-dependent measurements in the study of biological processes.
@@ Acknowledgements.
The authors are grateful for the support of the Air Force Office of Scientific Research (Award No. FA-9550-20-1-0366) and the Robert A. Welch Foundation (A-1943-20210327), National Science Foundation (Award No.2426699). The final measurements for the SNR were done in the quantum sensing lab at the University of Tennessee at Chattanooga (UTC). We wish to acknowledge the technical help provided by Charles Wallace at Texas AM University.Appendix A: Transfer-Matrix method
@@
Appendix B: Four-Wave Mixing process
The Four Wave Mixing process crucial in generating the TMBSS takes place inside a 2.5 cm long cylindrical vapor cell containing Rb-85 atomic ensemble. The atomic vapor serves as a non-linear medium where the line corresponding to the transition including the hyperfine structure of the Rb85 atom constitutes a double- energy level configuration as shown in Fig. 3(b).
Appendix C: Sensor assembly
@@
Appendix D: Shot Noise Limitation
We verified the coherent source for its shot-noise limit. The two coherent beams were directed in the two detectors A and B replicating the arrangement for TMBSS source. The verification was done using the oscilloscope for voltage readings and the spectrum analyzer was used to record the noise power around the working frequency of 2MHz. As shown in Fig. 11, units for y-axis were converted from the dBm units as displayed on the spectrum analyzer to and plotted against the total voltage . The difference in the scale of the units can be ascribed to using two different instruments as oscilloscope with impedance while the spectrum analyzer with . The linear relation between the mean of the signal and its variance confirms that our study was done in a shot noise limited domain.
Appendix E: SPR calibration curve
We perform the SPR Calibration under the exact instrumental conditions used for protein adsorption measurements by recording the reflectivity of the sensor against the external scanning angle for BSA (deionized water) and BSA solutions as shown in Fig. 12 using the coherent source. The scale on X-axis is calibrated such that the resonance angle for deionized water is at zero.
We then plot the resonance angle against each concentration as shown in Fig. 13. The plot shows the linearity as expected for a calibration curve. We compare this analysis with the theoretical expectations derived using the information provided in [42] about the refractive indices of various concentrations of the BSA solution. We map these refractive indices to the angle of resonance calculated using the Transfer Matrix Method (TMM) as explained in appendix A and the schematics section in this paper. We expect to see an approximate 2-degree shift as the percentage concentration increases by as confirmed in Fig. 13.